# Minimal Proper-time in Quantum Field Theory

### v1

@@ 1 Introduction
The key conceptual contribution of this work is the interpretation of the fundamental Planck scale as the inverse of the minimum proper time $\tau_{\min}>0$, which is automatically the same in all inertial frames. This may be seen as the other side of the medal of the minimal length scenario, often proposed in the literature [36, 25, 32, 38, 40, 31, 12, 13]. One motivation for considering a minimal length in flat space comes from general arguments from black hole physics and quantum mechanics [33, 6].
@@
The rest of this paper is organized as follows. In sec. 2, we highlight the main feature of the seminal work of Nambu [39]. In sec. 3, we recall the Schrödinger representation in QFT, specifically for a scalar real field, since it is not commonly used. Then, the succeeding section, 4, merges the concepts in the previous sections, namely, importing Nambu’s approach into QFT. With this formalism at hand, we introduce a minimal proper time, in the main section 5, as a fundamental physical scale. All the consequences of this scale are examined in sec. 6. Finally, we offer our conclusions in sec. 7. The article is completed by some appendices containing technical details: in the appendix A, we summarize the inner products in several formalism; in B, we elaborate on the free-particle physical picture in the light of the proper-time formalisms; in C, we revisit renormalization group in the proper-time language.
@@ 2 Overview on Nambu’s Proper-Time Quantum Mechanics
The eigenfunction with eigenvalue $\lambda=0$ is given by
$\psi^{R}(x)=\int_{0}^{\infty}d\tau\,\psi(x,\tau)=\int_{0}^{\infty}d\tau\,\psi(x)e^{-i\,\lambda\,\tau}=\delta(\lambda)\psi(x)\,$, (6)
where $\delta$ is the Dirac delta function. Thus, the physical states are constrained by the condition:
@@ 4 Nambu’s approach in QFT
One can obtain the usual $t$-independent Schrödinger equation for the relativistic Hamiltonian by writing,
$\Psi_{\lambda}[\phi,t]=\Psi_{\lambda}[\phi]\,e^{-i\,E\,\,t}$ (31)
to obtain
$\hat{H}\Psi_{\lambda}[\phi]$ $=(E+\lambda)\Psi_{\lambda}[\phi]$
$(\sqrt{\vec{p}^{2}+m^{2}})\Psi_{\lambda}[\phi]$ $=(E+\lambda)\Psi_{\lambda}[\phi]$ (32)
with $\omega_{\vec{p}}=\sqrt{\vec{p}^{2}+m^{2}}$. Applying $\hat{H}$ one more time to the equation above, one gets,
$(\vec{p}^{2}+m^{2})\Psi_{\lambda}[\phi]$ $=(E+\lambda)^{2}\Psi_{\lambda}[\phi]$
$\omega_{\vec{p}}^{2}\Psi_{\lambda}[\phi]$ $=(E^{2}+\lambda^{2}+2E\lambda)\Psi_{\lambda}[\phi]$
$(\lambda^{2}+2E\lambda+E^{2}-\omega_{\vec{p}}^{2})\Psi_{\lambda}[\phi]$ $=0\,$, (33)
When $E^{2}=\omega_{\vec{p}}^{2}$ (on-shell states), the eigenvalue equation has two solutions $\lambda=-2\omega_{\vec{p}}$ and $\lambda=0$. The solution $\lambda=0$ corresponds to states satisfying the Schrödinger equation, and the negative eigenvalue we regard as on-shell states not satisfying the Schrödinger equation.
@@
Integrating over $\tau$ the wave functional $\Psi[\phi,t,\tau]$ produces a wave functional satisfying the standard Schrödinger equation, namely
$\Psi^{R}[\phi,t]$ $\propto\int_{0}^{\infty}d\tau\,\Psi[\phi,t,\tau]$
$=\Psi_{\lambda}[\phi,t]\int_{0}^{\infty}d\tau\,e^{-i\lambda\tau}$
$\propto\delta(\lambda)\,\Psi_{\lambda}[\phi,t].$ (36)
@@ 4.1 Two-Point Function and Propagator in the Nambu-like formalism
which, integrated over $\tau$, gives,
$\int_{0}^{\infty}d\tau\,\langle 0,t,\tau|\hat{\phi}(\vec{x}^{\prime})\hat{\phi}(\vec{x})|0,0,0\rangle=\delta(\lambda)\int\mathcal{D}\phi\,\phi(\vec{x})\,\phi(\vec{x}^{\prime})\,\Psi_{0}^{*}[\phi,t]\Psi_{0}[\phi,0].$ (42)
Now recall that $\delta(\lambda)$ – i.e., $\lambda=0$ – selects the standard Schrödinger functional equation, thus leading to the standard propagator. Then, going to energy-momentum space, it must be
@@
with $\lambda_{p}=p_{0}-\omega_{\vec{p}}$, $\lambda_{a}=p_{0}+\omega_{\vec{p}}$, which are precisely the expressions found in (34) and (35) with $p_{0}=E$. Integrating over $\tau$ one recovers the standard Feynman propagator
$i\int_{0}^{\infty}d\tau\,\tau\,\int_{0}^{1}d\rho\,e^{-i\rho\lambda_{p}\tau-i(1-\rho)\lambda_{a}\tau}=\frac{i}{p^{2}-m^{2}}.$ (44)
@@
The expressions (44) and (45) are equivalent integral representations of the Feynman propagator. In Schwinger’s approach, the integral involves the proper time squared, whose generator is the four-momentum $p^{2}$, and we work with the proper time, whose generators are the eigenvalues of $\lambda$. In contrast, the integration limits of (44) and (45) are in correspondence, which will be relevant later. We elaborate on the interpretation of the proper-time representation of the Feynman propagator in Appendix B.
Once an interaction is turned on,
$H_{\text{int}}=\int d^{3}x\,\alpha\,\phi(\vec{x})^{4},$ (46)
the coupling $\alpha$ evolves according to renormalization group equations (RGEs). Usually, renormalization involves introducing a reference energy scale $\mu$ to absorb the cutoff dependence in loop integrals, and the RGEs describe how the coupling $\alpha(\mu)$ evolves with $\mu$ – see Appendix C.
@@ 5.1 Relaxing the constraint and exotic states
Consider the generalized functional Schrödinger equation with a parameter $\lambda$,
@@
$\Re\left[\int_{0}^{\infty}d\tau\,e^{-i\lambda\tau}\right]=\pi\delta(\lambda).$ (54)
Now, we introduce a minimal proper-time cutoff $\tau_{\min}>0$ and split the integral as:
$\int_{0}^{\infty}e^{-i\lambda\tau}d\tau=\int_{0}^{\tau_{\min}}e^{-i\lambda\tau}d\tau+\int_{\tau_{\min}}^{\infty}e^{-i\lambda\tau}d\tau.$ (55)
@@
$\Re\left[\int_{\tau_{\min}}^{\infty}e^{-i\lambda\tau}d\tau\right]=\pi\delta(\lambda)-\Re\left[\int_{0}^{\tau_{\min}}e^{-i\lambda\tau}d\tau\right].$ (56)
@@
$\Re\left[\cdot\right]=\frac{\sin(\lambda\tau_{min})}{\lambda}=\tau_{\min}+\mathcal{O}(\tau_{\min}^{3}),$ (57)
@@
$\Re\left[\int_{\tau_{\min}}^{\infty}e^{-i\lambda\tau}d\tau\right]=\pi\delta(\lambda)-\tau_{\min}+\mathcal{O}(\tau_{\min}^{3}),$ (58)
@@ 5.2 Superposition of Eigenstates
where the subscript $R$ denotes ’Real-world’, and with normalization $f(0)=1$, such that when $\tau_{min}\rightarrow 0$ one has the original constrained $\Psi_{(\lambda=0)}$. Similarly, when one has just $f(\lambda)=\delta(\lambda)$ one recovers $\Psi^{R}=\Psi_{(\lambda=0)}$.
@@ 6.3 Effective Planck constant
We observe that the theory contains a degree of freedom that can be exploited to clarify the meaning of $\hbar_{\text{eff}}(t)$. The key point is that the Nambu-like constraint can be imposed with a Dirac delta function having arbitrary normalization. Regardless of any prefactor in front of $\delta(\lambda)$ in (37), the constraint a la Nambu still selects $\lambda=0$. Similarly, in the presence of a soft violation due to $\tau_{\min}>0$, the function $f(\lambda)$ can be normalized with an arbitrary constant. Thus it reflects on $\tilde{f}(t)$ as a rescaling: $\tilde{f}(t)\rightarrow N\tilde{f}(t)$, where NN denotes the normalization factor. As a result, equation (84) becomes:
@@ 6.4 Effect on the running coupling behavior at high energies
At large $p_{E}^{2}$, the exponential factor $e^{-s_{\min}p_{E}^{2}}$ strongly suppresses UV contributions, effectively taming divergences and improving the ultraviolet behavior of the theory. Remaining in $\phi^{4}$-model, this mechanism ensures that loop integrals become finite and suggests the absence of a Landau pole, as in [34]. The reason is that the running of the coupling is no longer log-dominated at high energy, as in the standard case. Specifically, it is easy to work out the behavior of the four-point function, in the relevant regimes, using the propagator in (86):
$\alpha(\mu)\approx\alpha(\mu_{0})\left(1+\log(\frac{\mu}{\mu_{0}})-s_{min}\mu^{2}\right)\hskip 15.00002pt\mu\ll s_{min}^{-1/2}=\tau_{min}^{-1}$
@@
$\alpha(\mu)\approx\alpha(s_{min}^{-1/2})\left(1-\frac{e^{-2s_{min}\mu^{2}}}{4s_{min}\mu^{2}}\right)\hskip 36.0001pt\mu\gtrsim s_{min}^{-1/2}=\tau_{min}^{-1}.$ (87)
@@
The first line reproduces the standard, log behavior, as it must be. The second one shows that the coupling rapidly reaches a plateau in the UV, rendering $\phi^{4}$-model asymptotically safe. It effectively becomes scale invariant at very high energy.
A couple of comments are in order. First, the loops in the deep UV, for $E>\tau_{min}^{-1}$ are a sort of extrapolation, in our setting. In fact, the progressive vanishing of $\hbar_{eff}$ at UV in (85) suppresses the quantum correction, thus the loops. Consistently, the exponential suppression of the loops in deep UV from (86) gives an effective description of this phenomenon. Heuristically, the (UV) scale-invariance may be seen as a symptom of an underlying deterministic reality.
@@
Second, the regular UV behavior can be naturally interpreted as a dimensional reduction – see [18] for a review. Specifically, the term $\exp\left[-s(p_{E}^{2}+m^{2})\right]$ works as an entire function similarly to [45, 34, 14, 35] – an evolving dimension is discussed in [3, 19, 4], and the non-integer effective dimension can then be interpreted as fractal spacetime [38, 10, 15, 16, 17]. The underlying idea is that the effective dimension, different from the topological one ($D=4$), may run with the energy and lower in the UV. The UV behavior of the Green function is the one to measure the effective dimensionality [35]. Similarly, one can write a modified heat kernel, such that the corresponding spectral dimension is variable, vanishing at UV [37, 22].

### v2

@@ 1 Introduction
The key conceptual contribution of this work is the interpretation of the fundamental Planck scale as the inverse of the minimum proper time $\tau_{\min}>0$, which is automatically the same in all inertial frames. This may be seen as the other side of the medal of the minimal length scenario, often proposed in the literature [41, 28, 36, 46, 43, 45, 35, 15, 16]. One motivation for considering a minimal length in flat space comes from general arguments from black hole physics and quantum mechanics [37, 7].
@@
The rest of this paper is organized as follows. In sec. 2, we highlight the main feature of the seminal work of Nambu [44]. In sec. 3, we recall the Schrödinger representation in QFT, specifically for a scalar real field, since it is not commonly used. Then, the succeeding section, 4, merges the concepts in the previous sections, namely, importing Nambu’s approach into QFT. With this formalism at hand, we introduce a minimal proper time, in the main section 5, as a fundamental physical scale. All the consequences of this scale are examined in sec. 6. Finally, we offer our conclusions in sec. 7. The article is completed by some appendices containing technical details: in the appendix A, we summarize the inner products in several formalism; in C, we elaborate on the free-particle physical picture in the light of the proper-time formalisms; in D, we revisit renormalization group in the proper-time language.
@@ 2 Overview on Nambu’s Proper-Time Quantum Mechanics
The eigenfunction associated with the physical $\lambda=0$ sector can be extracted by integrating over the proper time $\tau$. Mathematically, this half-line integration yields a complex distribution,
$\int_{0}^{\infty}d\tau\,\psi(x,\tau)=\int_{0}^{\infty}d\tau\,\psi(x)e^{-i\,\lambda\,\tau}=\left[\pi\delta(\lambda)-i\mathcal{P}\left(\frac{1}{\lambda}\right)\right]\psi(x)\,$, (6)
where $\delta$ is the Dirac delta function and $\mathcal{P}$ denotes the Cauchy principal value. As we will discuss in more detail later (see Eq. 53), it is the real part of this integration that properly enforces the $\lambda=0$ constraint, projecting the extended state onto the standard physical wave function. We can thus formally define the recovered physical state $\psi^{R}(x)$ by isolating this delta-function contribution:
$\psi^{R}(x)\propto\text{Re}\left(\int_{0}^{\infty}d\tau\,\psi(x,\tau)\right)=\pi\delta(\lambda)\psi(x)\,.$ (7)
Thus, the physical states are constrained by the condition:
@@ 4 Nambu’s approach in QFT
Let us clarify how the extended proper-time dynamics translates into a shifted energy eigenvalue for the functional Hamiltonian. Starting from the generalized Nambu-like equation in the $\lambda$-sector, we have:
$(\hat{H}-i\partial_{t})\Psi_{\lambda}[\phi,t]=\lambda\Psi_{\lambda}[\phi,t]\,$, (32)
which can be rearranged as $\hat{H}\Psi_{\lambda}[\phi,t]=i\partial_{t}\Psi_{\lambda}[\phi,t]+\lambda\Psi_{\lambda}[\phi,t]$. We can map this into the standard functional Schrödinger equation by performing a phase redefinition. Let us define a new functional $\bar{\Psi}_{\lambda}[\phi,t]$ such that
$\Psi_{\lambda}[\phi,t]=e^{i\lambda t}\bar{\Psi}_{\lambda}[\phi,t]\,.$ (33)
Substituting this into the equation above, the $\lambda$ terms cancel out, leaving exactly the standard functional Schrödinger equation for the redefined state:
$\hat{H}\bar{\Psi}_{\lambda}[\phi,t]=i\partial_{t}\bar{\Psi}_{\lambda}[\phi,t]\,.$ (34)
If we now consider stationary states for the standard equation, we have $\bar{\Psi}_{\lambda}[\phi,t]=\Psi_{\lambda}[\phi]e^{-i\bar{E}t}$, where $\bar{E}$ is the eigenvalue of $\hat{H}$. Consequently, the original wave functional takes the form
$\Psi_{\lambda}[\phi,t]=\Psi_{\lambda}[\phi]e^{-i(\bar{E}-\lambda)t}\,.$ (35)
By identifying the physical, observed energy as $E\equiv\bar{E}-\lambda$ (so that $\bar{E}=E+\lambda$), we write $\Psi_{\lambda}[\phi,t]=\Psi_{\lambda}[\phi]e^{-iEt}$, which directly yields the shifted eigenvalue equation for the functional Hamiltonian:
$\hat{H}\Psi_{\lambda}[\phi]=(E+\lambda)\Psi_{\lambda}[\phi]\,.$ (36)
At this stage, $\Psi_{\lambda}[\phi]$ is a general wave functional describing an arbitrary field configuration. To address the particle spectrum and make contact with the standard momentum-space dispersion relations, we evaluate this equation for a single-particle eigenfunctional with momentum $\vec{p}$. In the functional Schrödinger representation, such a state is explicitly given by  [^2] See Eq. 10.30 in Ref. [32].
$\Psi_{\lambda,\vec{p}}[\phi]\equiv\tilde{\phi}(\vec{p})\Psi_{0}[\phi]=\mathcal{N}\left(\int d^{3}x\,e^{-i\vec{p}\cdot\vec{x}}\,\phi(\vec{x})\right)\Psi_{0}[\phi]\,$, (37)
where $\Psi_{0}[\phi]$ is the vacuum wave functional and $\mathcal{N}$ is a normalization constant, and it is of the form [32]
$\Psi_{0}[\phi]=\eta\,e^{-\int\,d^{3}x\,d^{3}y\,\phi(\vec{x})\,g(\vec{x},\vec{y})\,\phi(\vec{y})}$ (38)
where $\eta$ is a normalization constant whose specific form is not relevant for our discussion and
$g(\vec{x},\vec{y})=\int\frac{d^{3}k}{(2\pi)^{3}}\tilde{g}(\vec{k})e^{i\vec{k}\cdot(\vec{x}-\vec{y})}$ (39)
with
$\tilde{g}^{2}(\vec{k})=\frac{1}{4}\left(\vec{k}^{2}+m^{2}\right)\,.$ (40)
For the specific state $\Psi_{\lambda,\vec{p}}[\phi]$, the free functional Hamiltonian in (17) yields the eigenvalue
$\omega_{\vec{p}}=\sqrt{\vec{p}^{2}+m^{2}},$
so that
$\hat{H}\Psi_{\lambda,\vec{p}}[\phi]=\omega_{\vec{p}}\,\Psi_{\lambda,\vec{p}}[\phi].$ (41)
The eigenvalue equation then becomes
$\omega_{\vec{p}}\,\Psi_{\lambda,\vec{p}}[\phi]=(E+\lambda)\Psi_{\lambda,\vec{p}}[\phi]\,.$ (42)
Applying $\hat{H}$ one more time to the equation above (i.e., acting with the squared energy operator), one gets
$(\vec{p}^{2}+m^{2})\Psi_{\lambda,\vec{p}}[\phi]$ $=(E+\lambda)^{2}\Psi_{\lambda,\vec{p}}[\phi]$
$\omega_{\vec{p}}^{2}\Psi_{\lambda,\vec{p}}[\phi]$ $=(E^{2}+\lambda^{2}+2E\lambda)\Psi_{\lambda,\vec{p}}[\phi]$
$(\lambda^{2}+2E\lambda+E^{2}-\omega_{\vec{p}}^{2})\Psi_{\lambda,\vec{p}}[\phi]$ $=0\,.$ (43)
When $E^{2}=\omega_{\vec{p}}^{2}$ (on-shell states), the eigenvalue equation has two solutions $\lambda=-2\omega_{\vec{p}}$ and $\lambda=0$. The solution $\lambda=0$ corresponds to states satisfying the standard Schrödinger equation, and the negative eigenvalue we regard as on-shell states not satisfying the Schrödinger equation.
@@
Integrating over $\tau$ the wave functional $\Psi[\phi,t,\tau]$ produces a wave functional satisfying the standard Schrödinger equation, namely, using (7),
$\Psi^{R}[\phi,t]\propto\text{Re}\left(\int_{0}^{\infty}d\tau\,\Psi[\phi,t,\tau]\right)\ \propto\delta(\lambda)\,\Psi_{\lambda}[\phi,t].$ (46)
@@ 4.1 Two-Point Function and Propagator in the Nambu-like formalism
which, integrated over $\tau$ and extracting the real part to enforce the constraint as discussed for the $\lambda=0$ sector, gives
$\text{Re}\left(\int_{0}^{\infty}d\tau\,\langle 0,t,\tau|\hat{\phi}(\vec{x}^{\prime})\hat{\phi}(\vec{x})|0,0,0\rangle\right)\propto\delta(\lambda)\int\mathcal{D}\phi\,\phi(\vec{x})\,\phi(\vec{x}^{\prime})\,\Psi_{0}^{*}[\phi,t]\Psi_{0}[\phi,0].$ (52)
Now recall that isolating $\delta(\lambda)$ – i.e., enforcing $\lambda=0$ – selects the standard Schrödinger functional equation, thus leading to the standard propagator. Then, going to energy-momentum space, one can construct the proper-time transition amplitude. To correctly reproduce the time-ordered causal structure of QFT, one must symmetrically include both the particle and antiparticle off-shell branches. This is achieved by introducing a mixing parameter $\rho\in[0,1]$, akin to a Feynman parameter, which linearly interpolates between the two solutions. The amplitude is then written as:
@@
with $\lambda_{p}=p_{0}-\omega_{\vec{p}}$ and $\lambda_{a}=p_{0}+\omega_{\vec{p}}$, which are precisely the expressions found in (44) and (45) with $p_{0}=E$. The parameter $\rho$ essentially ensures that both positive and negative frequency modes are on the same footing. Integrating over $\tau$ one recovers exactly the standard Feynman propagator
$i\int_{0}^{\infty}d\tau\,\tau\,\int_{0}^{1}d\rho\,e^{-i\rho\lambda_{p}\tau-i(1-\rho)\lambda_{a}\tau}=\frac{i}{p^{2}-m^{2}+i\epsilon}.$ (54)
(Note: we implicitly assume the standard $i\epsilon$ prescription to ensure the convergence of the $\tau$ integration at infinity, as detailed in Appendix B).
@@
The expressions (54) and (55) are equivalent integral representations of the Feynman propagator. In Schwinger’s approach, the integral involves the proper time squared, whose generator is the four-momentum $p^{2}$, and we work with the proper time, whose generators are the eigenvalues of $\lambda$. In contrast, the integration limits of (54) and (55) are in correspondence, which will be relevant later. We elaborate on the interpretation of the proper-time representation of the Feynman propagator in Appendix C.
We have analyzed the free model in full detail, which serves as the basis for quantization aiming to treat the interaction as a perturbation. The canonical quantization procedure is performed on the free Lagrangian, whose exact solvability allows for a well-defined Fock space construction. Interactions are subsequently introduced as perturbations built upon this free, quantized structure. This methodology follows the conventional formulation of canonical quantization in QFT. Once an interaction is turned on,
$H_{\text{int}}=\int d^{3}x\,\frac{\alpha}{4!}\,\phi(\vec{x})^{4},$ (56)
the coupling $\alpha$ evolves according to renormalization group equations (RGEs). Usually, renormalization involves introducing a reference energy scale $\mu$ to absorb the cutoff dependence in loop integrals, and the RGEs describe how the coupling $\alpha(\mu)$ evolves with $\mu$ – see Appendix D.
@@ 5.1 Relaxing the constraint and exotic states
Naively, the introduction of a minimal proper-time $\tau_{min}$ makes the right-hand side of Eq. (27) ill-defined. To solve this issue, one can resort to an analogy with the approach of [40]. The functional $\Psi[\phi,t,\tau]$ can be made differentiable by formally introducing a $\tau_{min}$ dependence in the functional $\Psi[\phi,t,\tau]\to\Psi[\phi,t,\tau\,|\,\tau_{min}]$ via a smoothing function. In other words, we regularize $\Psi$:
$\Psi_{reg}[\phi,t,\tau]=\int_{0}^{\infty}d\tau^{\prime}S(\tau-\tau^{\prime}|\tau_{min})\Psi[\phi,t,\tau^{\prime}]\,$, (58)
where the subscript ’regreg’ denotes the regularized functional, and $S$ is a smoothing kernel. For the sake of notation, in what follows we will omit the subscript ’regreg’ and the explicit $\tau_{min}$ dependence. Changes in the scale $\tau_{min}$ are, as usual, governed by the renormalization group equation, as discussed in appendix D.
Once the functional $\Psi[\phi,t,\tau]$ is properly regularized, one can consider the generalized functional Schrödinger equation with a parameter $\lambda$,
@@
$\text{Re}\left[\int_{0}^{\infty}d\tau\,e^{-i\lambda\tau}\right]=\pi\delta(\lambda).$ (65)
To promote the introduction of a minimal proper time $\tau_{\min}$ from a mere regularization device to a structural assumption of the formalism, we postulate that the proper-time representation of the theory is not operationally defined below a fundamental lower bound $\tau_{\min}>0$. In other words, phenomena associated with scales shorter than $\tau_{\min}$ are assumed to lie outside the domain of validity of the effective description.
Under this assumption, the strict projection onto the physical sector obtained by integrating down to $\tau=0$ — which yields an exact $\delta(\lambda)$ constraint — is interpreted as an idealized limit. At finite resolution, the lower bound $\tau_{\min}$ prevents this exact projection, replacing it with a smeared constraint in which modes with $\lambda\neq 0$ are not completely eliminated.
Accordingly, the proper-time integral defining the physical sector must explicitly reflect this minimal scale. We therefore decompose
$\int_{0}^{\infty}e^{-i\lambda\tau}\,d\tau=\int_{0}^{\tau_{\min}}e^{-i\lambda\tau}\,d\tau+\int_{\tau_{\min}}^{\infty}e^{-i\lambda\tau}\,d\tau\,$, (66)
and interpret the contribution from the interval $0,\tau_{\min})$ as lying beyond the regime of validity of the effective theory.
@@
$\text{Re}\left[\int_{\tau_{\min}}^{\infty}e^{-i\lambda\tau}d\tau\right]=\pi\delta(\lambda)-\text{Re}\left[\int_{0}^{\tau_{\min}}e^{-i\lambda\tau}d\tau\right].$ (67)
@@
$\text{Re}\left[\cdot\right]=\frac{\sin(\lambda\tau_{min})}{\lambda}=\tau_{\min}+\mathcal{O}(\tau_{\min}^{3}),$ (68)
@@
$\text{Re}\left[\int_{\tau_{\min}}^{\infty}e^{-i\lambda\tau}d\tau\right]=\pi\delta(\lambda)-\tau_{\min}+\mathcal{O}(\tau_{\min}^{3}),$ (69)
@@ 5.2 Superposition of Eigenstates
where the subscript $R$ denotes ’Real-world’, such that when $\tau_{min}\rightarrow 0$ one has the original constrained $\Psi_{(\lambda=0)}$. Similarly, in opposite logical direction, when one has simply $f(\lambda)=\delta(\lambda)$ one recovers $\Psi^{R}=\Psi_{(\lambda=0)}$ and $\tau_{min}=0$. Notice that this $\tau_{min}$ dependence in $f(\lambda)$ emanates from the regularization of the functional, namely, (58).
@@ 6.3 Effective Planck constant
We observe that the theory contains a degree of freedom that can be exploited to clarify the meaning of $\hbar_{\text{eff}}(t)$. The key point is that the Nambu-like constraint can be imposed with a Dirac delta function having arbitrary normalization. Regardless of any prefactor in front of $\delta(\lambda)$ in (47), the constraint a la Nambu still selects $\lambda=0$. Similarly, in the presence of a soft violation due to $\tau_{\min}>0$, the function $f(\lambda)$ can be normalized with an arbitrary constant. Thus it reflects on $\tilde{f}(t)$ as a rescaling: $\tilde{f}(t)\rightarrow N\tilde{f}(t)$, where $N$ denotes the normalization factor. As a result, equation (95) becomes:
@@ 6.4 Effect on the running coupling behavior at high energies
At large $p_{E}^{2}$, the exponential factor $e^{-s_{\min}p_{E}^{2}}$ strongly suppresses UV contributions, effectively taming divergences and improving the ultraviolet behavior of the theory. Remaining in $\phi^{4}$-model, this mechanism ensures that loop integrals become finite and suggests the absence of a Landau pole, as in [38]. The reason is that the running of the coupling is no longer log-dominated at high energy, as in the standard case. Specifically, it is easy to work out the behavior of the four-point function $\alpha/4!\,\phi(x)^{4}$ model, in the relevant regimes, using the propagator in (97):
$\alpha(\mu)\approx\alpha(\mu_{0})\left(1+\alpha(\mu_{0})\,\beta_{0}\log(\frac{\mu}{\mu_{0}})-\alpha(\mu_{0})s_{min}\mu^{2}\right)\hskip 15.00002pt\mu\ll s_{min}^{-1/2}=\tau_{min}^{-1}$
@@
$\alpha(\mu)\approx\alpha(s_{min}^{-1/2})\left(1-\beta_{0}\,\alpha(s_{min}^{-1/2})\frac{e^{-2s_{min}\mu^{2}}}{4s_{min}\mu^{2}}\right)\hskip 36.0001pt\mu\gtrsim s_{min}^{-1/2}=\tau_{min}^{-1},$ (98)
@@
where $\beta_{0}=3/(16\pi^{4})$. The first line reproduces the standard, log behavior, as it must be. The second one shows that the coupling rapidly reaches a plateau in the UV, rendering $\phi^{4}$-model asymptotically safe. It effectively becomes scale invariant at very high energy.
Figure 1: Qualitative comparison of the standard $\phi^{4}$ running coupling and its version with a minimal proper time. The two coincide in the infrared, while the new UV effects become relevant above the scale $\tau_{min}$. In contrast to the standard prediction, which develops a Landau pole, the running in the new framework remains finite in the ultraviolet.
Some implications are now highlighted.
  1. 1.
    A conceptual obstacle for perturbative QFT is the presence of the UV renormalons which destroy the Borel resummability of the perturbative series [1]. The renormalons arise from the logarithm behavior of the running coupling in deep UV – this is particularly manifest from the resurgent approach in [12, 39] – causing the appearance of factorial divergent contribution in the power series. In the present framework, the modification due to $\tau_{min}$ in (98) modifies the standard logarithm behavior as also illustrated in Fig. 1, then eliminating the source of the UV renormalon ambiguities[38]. This is a nontrivial result: the absence of renormalons indicates that the perturbative expansion remains well-controlled. Combined with the automatic and guaranteed asymptotic safety of the theory, this ensures that the renormalization procedure is internally consistent and free from pathological ambiguities at arbitrarily high energies.
    Additionally, the absence of logarithmic running of the coupling in the ultraviolet (UV) regime invalidates ’t Hooft’s argument for the so-called horn-shaped singularities in the Green functions [1]. As shown in [38], without logarithmic behavior the perturbative expansion does not exhibit super-exponential growth in Borel space. This removes a potential source of pathological UV behavior in perturbation theory and significantly reinforces the theoretical consistency of the framework at short distances, while preserving all standard QFT predictions at low energies.
    These considerations further motivate our strategy: we first analyze the free-field theory in detail and subsequently introduce the interaction term in (56), to be treated perturbatively.
  2. 2.
    In the deep UV regime, $E>\tau_{\min}^{-1}$, loop effects should be regarded as a formal extrapolation within our framework. Indeed, the progressive vanishing of $\hbar_{\text{eff}}$ in the UV, see (96), suppresses quantum corrections and therefore loop contributions. This behavior is consistently reflected in the exponential damping of the propagator in (97), which provides an effective description of the same mechanism. As a result, the $\beta$-function asymptotically vanishes and exact scale invariance is recovered in the ultraviolet.
  3. 3.
@@
    The regular UV behavior can be naturally interpreted as a dimensional reduction – see [21] for a review. Specifically, the term $\exp\left[-s(p_{E}^{2}+m^{2})\right]$ works as an entire function similarly to [51, 38, 17, 40] – an evolving dimension is discussed in [4, 22, 5], and the non-integer effective dimension can then be interpreted as fractal spacetime [43, 11, 18, 19, 20]. The underlying idea is that the effective dimension, different from the topological one ($D=4$), may run with the energy and lower in the UV. The UV behavior of the Green function is the one to measure the effective dimensionality [40]. Similarly, one can write a modified heat kernel, such that the corresponding spectral dimension is variable, vanishing at UV [42, 25].